Operations Report – November 13-16th

MDRS Operations Report 13-16-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Tuesday):

Average 50.85

Minimum 47.50

Maximum 58.0

Notes on solar: Cloudy until this afternoon but SOC was 99% at sunset

Propane Reading, station tank – not checked %

Propane Reading, director tank— not checked %

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): not checked gallons

Water Meter: not checked units

Water (static tank) – not checked gallons

Static to Loft Pump used – no

Water in GreenHab – 20 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Notes on rovers: Unplugged due to the cloudy weather today

ATV’s Used: (Honda, 350.1, 350.2, 350.3): no

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, to town to dump trash

CrewCar used and why, where? No

Luna used and why, where? Yes, to Bicknell

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal.

Summary of suits and radios: Nothing to report.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: nothing to report

Summary of Hab operations: Did a complete propane check on Monday and still no evidence of any propane problems. Heaters are off, but the hot water heater, which was what the crew reported as leaking is working nominally.

Summary of Outpost operations: My shower stopped working. No idea why.

Summary of GreenHab operations: Fertilized the plants. Between this and the increase in temperature, I have seen an increase in growth. I need to get the supplemental lighting set up but worry about the increase in nighttime power use. I am looking into switching what we currently have with LED lights.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at report.